OpenStudy (anonymous):

Given that the acceleration of gravity at the surface of Mars is 0.38 of what it is on Earth, and that Mars' radius is 3400 km, determine the mass of Mars. please show me how to do this!!!

OpenStudy (schrodingers_cat):

Remember the the acceleration felt by gravity on an object from another object would be given by a = GM/(r^2) So, if mars' gravity is .38 of that of earths it would be 9.8 x .38 = 3.724 m/s^2 Then you can use the equation above to get mars' mass M = 3.724((3400 x1000)^2)/G = 6.45 x 10^23 kg Hope this helps :)

OpenStudy (anonymous):

thanks so much